# CORE-3D: Context-aware Open-vocabulary Retrieval by Embeddings in 3D

### v2

@@ Abstract
3D scene understanding is fundamental for embodied AI and robotics, supporting reliable perception for interaction and navigation. Recent approaches achieve zero-shot, open-vocabulary 3D semantic mapping by assigning embedding vectors to 2D class-agnostic masks generated via vision-language models (VLMs) and projecting these into 3D. However, these methods often produce fragmented masks and inaccurate semantic assignments due to the direct use of raw masks, limiting their effectiveness in complex environments. To address this, we leverage SemanticSAM with progressive granularity refinement to generate more accurate and numerous object-level masks, mitigating the over-segmentation commonly observed in mask generation models such as vanilla SAM, and improving downstream 3D semantic segmentation. To further enhance semantic context, we employ a context-aware CLIP encoding strategy that integrates multiple contextual views of each mask using empirically determined weighting, providing much richer visual context. We evaluate our approach on multiple 3D scene understanding tasks, including 3D semantic segmentation and object retrieval from language queries, across several benchmark datasets. Experimental results demonstrate significant improvements over existing methods, highlighting the effectiveness of our approach.
@@ 1 Introduction
  • We enforce reliable multi-view semantic 3D map by merging overlapping predictions in 3D and filtering incomplete or spurious masks using geometric heuristics.
@@ 2.1 Foundation Models for Vision–Language Alignment
Large-scale vision–language models have become the cornerstone of open-vocabulary perception. Contrastive pretraining approaches such as CLIP (Radford et al., 2021) learn aligned image and text embeddings from massive web corpora, enabling zero-shot classification, retrieval, and multimodal reasoning without task-specific finetuning. Building on this paradigm, ALIGN (Jia et al., 2021) and Florence (Yuan et al., 2021) improved representation quality, while region-level extensions such as GLIP (Li et al., 2022b) and GroupViT (Xu et al., 2022) extended open-vocabulary capabilities to detection and segmentation. Subsequently, BLIP (Li et al., 2022a) and BLIP-2 (Li et al., 2023c) enhanced grounding and enabled integration with large language models. Also, OVSeg (Liang et al., 2023) addresses the CLIP bottleneck in two-stage segmentation by finetuning on masked regions and introducing mask prompt tuning. These models provide rich semantic embeddings that are widely adopted as the backbone for open-vocabulary 2D and 3D pipelines.
@@ 2.2 Foundation Models for Mask Generation
Complementing semantic embeddings, class-agnostic segmentation priors have shown remarkable generalization across diverse domains. MaskFormer (Cheng et al., 2021) unified semantic and instance segmentation by reformulating both tasks as per-pixel mask classification, while its successor Mask2Former (Cheng et al., 2022) introduced masked attention to achieve strong panoptic segmentation performance with improved efficiency. MaskDINO (Li et al., 2023b) further integrates detection and segmentation in a unified transformer, showing strong generalization to unseen categories. Also, FreeSeg (Qin et al., 2023) proposes a unified framework that handles semantic, instance, and panoptic segmentation through multi-task training and adaptive prompt learning. The Segment Anything Model (SAM) (Kirillov et al., 2023), trained on billions of masks, demonstrated that a single backbone can transfer across domains and serves as a universal prior for open-vocabulary pipelines. However, SAM often produces fragmented or incomplete masks in cluttered indoor scenes. SemanticSAM (Li et al., 2023a) alleviates this through progressive multi-granularity refinement, yielding more coherent object-level masks particularly beneficial for downstream tasks requiring consistent segmentation.
2.3 Open-Vocabulary 3D Scene Understanding
@@ 3.1 Mask Generation
To address this, we introduce a progressive refinement strategy. For each image $I_{t}$, we generate segmentations at an increasing sequence of granularity levels $\{g_{1},g_{2},\ldots,g_{K}\}$. At each step $k$, SemanticSAM produces a set of $N_{k}$ candidate masks, denoted as
@@
The final mask set for frame $t$ is then
$\mathcal{M}_{t}=\bigcup_{k=1}^{K}\hat{\mathcal{M}}_{t}^{(k)}.$
In addition, we apply two lightweight filtering steps to further improve mask quality: (i) removing very small or marginal masks whose area falls below a minimum threshold, and (ii) applying DBSCAN clustering (Ester et al., 1996) directly on the pixel regions of each mask to merge fragmented parts and suppress spurious subdivisions of the same object.
@@ 3.2 Context-Aware CLIP Embedding
Given the refined 2D masks $\mathcal{M}_{t}$, we next compute semantic embeddings for each object candidate. A direct approach would be to crop the mask region and feed it into CLIP (Radford et al., 2021). However, CLIP relies heavily on visual context, and isolated object crops often lead to ambiguous or incorrect embeddings (e.g., a cropped chair leg being misinterpreted as a stick). To mitigate this, we construct a set of complementary visual crops for each mask that balance object detail with surrounding scene context.
@@
With both 2D masks $\mathcal{M}_{t}$ and their associated embeddings $\mathbf{e}(m)$, we lift the masks into 3D to enforce multi-view consistency and refinement. Given depth maps $D_{t}$ and camera poses $\mathbf{T}_{t}$, each pixel $p=(u,v)$ belonging to a mask $m$ is back-projected into 3D: (Hartley & Zisserman, 2003)
$\mathbf{x}_{p}=\mathbf{T}_{t}\begin{bmatrix}D_{t}(p)\,K^{-1}\tilde{p}\\ 1\end{bmatrix}$,
where $K$ is the camera intrinsics matrix and $\tilde{p}=(u,v,1)^{\top}$ denotes the homogeneous pixel coordinate.
@@
When a single 2D mask spans multiple distinct objects that are spatially close in the image but separated in 3D space (e.g., a vase in front of a couch), we refine it by applying DBSCAN clustering (Ester et al., 1996) to $\mathcal{X}(m)$. This step separates the projected 3D points of $m$ into distinct clusters, each corresponding to a potential object. Each cluster is treated as a new candidate instance: we project it back to the image plane, generate the corresponding 2D mask, and re-apply our multi-crop CLIP embedding procedure. This ensures that each physically distinct object obtains its own semantic descriptor, even if they were originally fused into a single 2D segmentation.
After applying merging and splitting, we obtain the final set of refined 3D masks:
@@ 3.4 Object Retrieval
where $m$ denotes the name and attributes of the main object , $\mathcal{R}$ is a set of referenced object names and $\Omega$ encodes orientation constraints (e.g., “front of the cabinet”). A lightweight LLM extractor produces $\Pi(q)$ deterministically.
Candidate mining. For each object name $x\in\{m\}\cup\mathcal{R}$, we compute CLIP similarity using pre-computed object embeddings and retain the top-$K$ matches. Each match corresponds to a 3D point cluster. To remove duplicates, we voxelize the clusters and discard clusters that overlap more than a fix threshold with a larger neighbor.
View selection and VLM verification. Each candidate cluster is then projected into the set of RGB frames. We select the frame that maximizes 3D–2D overlap while penalizing occlusion by other candidate clusters. A VLM is prompted with the bounding box region and asked a binary question to determine whether the object is present. Only candidates passing this check are retained.
Orientation grounding. If $\Omega$ specifies an orientation, we collect canonical views of each candidate at discretized yaw bins. These views are tiled into a numbered grid, and a VLM is asked to select the index corresponding to the orientation token (e.g., front). The chosen index is mapped back to a yaw angle and stored with the candidate.
Final reasoning. The remaining candidates for $m$, together with centroids of related objects and any orientation cues, are passed to an LLM. The LLM receives the original query and structured scene geometry and outputs the index of the final prediction.
@@ 4 Experiments
We evaluate our framework on two tasks: (i) 3D open-vocabulary semantic segmentation, where the goal is to assign category labels to 3D object instances without task-specific training, and (ii) natural-language object retrieval, where the goal is to localize objects in a 3D scene given free-form text queries that may contain relational and orientation constraints. All experiments are conducted on a workstation equipped with a single NVIDIA RTX4090.
@@
Datasets. We conduct experiments on two standard benchmarks: Replica (Straub et al., 2019) and ScanNet (Dai et al., 2017). Replica provides high-quality synthetic RGB-D scans of indoor environments with accurate ground-truth meshes and semantic annotations, while ScanNet consists of large-scale real-world RGB-D sequences with manually annotated 3D semantic and instance labels. Following prior work, we use eight Replica scenes: room0, room1, room2, office0, office1, office2, office3, and office4, and eight ScanNet scenes: 0011_00, 0030_00, 0046_00, 0086_00, 0222_00, 0378_00, 0389_00, and 0435_00. This subset selection ensures comparability with previous zero-shot methods. For text–image alignment, we use the Eva2 CLIP (Fang et al., 2023; Yang et al., 2023) vision-language model.
Labeling protocol. For open-vocabulary labeling, we use the standard CLIP paradigm (Radford et al., 2021): each refined 3D object mask is encoded into an embedding and compared against a set of text embeddings corresponding to candidate categories. The category with the highest similarity is assigned as the predicted label. This procedure enables zero-shot semantic segmentation without task-specific training, while our multi-view refinement makes the assignments more robust to occlusions and clutter.
Evaluation. We follow the evaluation protocol in prior work. For each predicted point cloud instance, we assign a semantic label by finding the nearest ground-truth instance (based on centroid distance) and transferring its label. For each scene, we restrict the text prompts to the classes present in the ground-truth annotations, formatted as “a photo of <class name>.” We report mean accuracy (mAcc), mean intersection-over-union (mIoU), and frequency-weighted mIoU (fmIoU).
Baselines. We compare against recent zero-shot approaches for 3D semantic segmentation: ConceptFusion (Jatavallabhula et al., 2023), ConceptGraphs (Gu et al., 2024), BBQ-CLIP (Linok et al., 2025), OpenMask3D (Takmaz et al., 2023), and HOV-SG (Werby et al., 2024). Note that HOV-SG reports results on a different subset of ScanNet, so we omit its numbers for fair comparison.
Results. Table 1 reports quantitative comparisons. Our method achieves the best performance across all three metrics, outperforming previous zero-shot methods. In particular, the improvements in mIoU and fwIoU highlight the benefit of our context-aware embeddings and multi-view 3D refinement, which produce more consistent object representations than 2D-based pipelines. We also provide qualitative examples in Fig. 4. Visual comparisons on Replica scenes illustrate that our method yields more accurate segmentation boundaries and finer recognition of challenging categories. Our method consistently detects objects across categories with higher fidelity than competing approaches.
@@ 4.2 Natural-Language Object Retrieval
Implementation details. We use the EVA02 CLIP backbone (Fang et al., 2023) for visual–text alignment, and qwen2.5-vl-32b-instruct as the vision–language model for multimodal encoding. Our first LLM, gpt-5-mini, is employed for object extraction, while the second LLM, openai-o4-mini, handles the final decision-making stage.
@@
Evaluation. Following prior work, we report grounding accuracy at two IoU thresholds: A@0.1 and A@0.25. A prediction is considered correct if the IoU between the predicted and ground-truth bounding box exceeds the threshold. Accuracy is reported overall as well as separately for the four difficulty subsets.
@@ 5 Conclusion
Author Contributions
M. Amin Mirzaei developed the 3D embeddings map generation framework and contributed to the 3D semantic segmentation experiments and literature review and also led the project. Pantea Amoie conducted the literature review, collaborated on the 3D semantic segmentation experiments, and wrote the entire manuscript. Ali Ekhterachian developed the object retrieval methodology, co-conducted the retrieval experiments, and revised the manuscript. Matin Mirzababaei co-developed the object retrieval methodology, conducted the retrieval experiments, and revised the manuscript.

### v3

@@ Abstract
3D scene understanding is fundamental for embodied AI and robotics, supporting reliable perception for interaction and navigation. Recent approaches achieve zero-shot, open-vocabulary 3D semantic mapping by assigning embedding vectors to 2D class-agnostic masks generated via vision-language models (VLMs) and projecting these into 3D. However, these methods often produce inaccurate semantic assignments due to the direct use of raw masks, limiting their effectiveness in complex environments. To address this, we leverage SemanticSAM with progressive granularity refinement to generate more accurate and numerous object-level masks. To further enhance semantic context, we employ a context-aware CLIP encoding strategy that integrates multiple contextual views of each mask, providing much richer visual context. We also propose a new VLM based framework to process complex 3D queries and retrieve target ojects based on 3D position relations between different objects. We evaluate our approach on multiple 3D scene understanding tasks, including 3D semantic segmentation and object retrieval from language queries, across several benchmark datasets. Experimental results demonstrate significant improvements over existing methods, highlighting the effectiveness of our approach.
@@ 1 Introduction
  • We enforce reliable multi-view semantic 3D map by merging overlapping predictions in 3D using geometric and semantic heuristics.
@@ 2.1 Foundation Models for Vision–Language Alignment
Large-scale vision–language models have become the cornerstone of open-vocabulary perception. Contrastive pretraining approaches such as CLIP (Radford et al., 2021) learn aligned image and text embeddings from massive web corpora, enabling zero-shot classification, retrieval, and multimodal reasoning without task-specific finetuning. Building on this paradigm, ALIGN (Jia et al., 2021) and Florence (Yuan et al., 2021) improved representation quality. Subsequently, OVSeg (Liang et al., 2023) addresses the CLIP bottleneck in two-stage segmentation by finetuning on masked regions and introducing mask prompt tuning. But even this pipeline can not fully resolve the problem of generating highly enriched embeddings for 2D masks.
@@ 2.2 Foundation Models for Mask Generation
Complementing semantic embeddings, class-agnostic segmentation priors have shown remarkable generalization across diverse domains. MaskFormer (Cheng et al., 2021) unified semantic and instance segmentation by reformulating both tasks as per-pixel mask classification, while its successor Mask2Former (Cheng et al., 2022) introduced masked attention to achieve strong panoptic segmentation performance with improved efficiency. MaskDINO (Li et al., 2023b) further integrates detection and segmentation in a unified transformer, showing strong generalization to unseen categories. The Segment Anything Model (SAM) (Kirillov et al., 2023), trained on billions of masks, demonstrated that a single backbone can transfer across domains and serves as a universal prior for open-vocabulary pipelines. However, SAM often produces fragmented or incomplete masks in cluttered indoor scenes. SemanticSAM (Li et al., 2023a) alleviates this by generating Segmnet masks with different granularity levels from semantic to instance and to parts.
2.3 Training-based 3D Scene Understanding
There are several methods that train a separate network for each scene to predict 3D semantic embeddings or labels for individual 3D points. These approaches typically extract embedding features from 2D image crops or masks of the scene and then assign them to the points. LERF (Kerr et al., 2023) utilizes the structure of NeRF (Mildenhall et al., 2020) models to assign not only colors and 3D positions, but also embedding vectors generated by the CLIP model to 3D points. While OpenNeRF (Engelmann et al., 2024) leverages OpenSeg (Ghiasi et al., 2022) feature maps to supervise the 3D feature maps.
Some models blend this idea with 3D Gaussian Splatting (Kerbl et al., 2023) structure. LangSplat (Qin et al., 2024) uses CLIP embeddings of SAM-generated masks to train an encoder–decoder architecture that produces low-dimensional semantic features for 3D Gaussian distributions. OpenGaussian (Wu et al., 2024), on the other hand, introduces an intra-mask smoothing loss and an inter-mask contrastive loss to enhance performance on 3D semantic segmentation tasks.
2.4 Zero-shot 3D Scene Understanding
@@ 3.1 Mask Generation
To address this, we introduce a progressive refinement strategy. For each image $I_{t}$, we generate segmentations at an increasing sequence of granularity levels $\{g_{1},g_{2},\ldots,g_{K}\}$. At each step $k$, SemanticSAM produces a set of candidate masks, but we only keep the set of masks $N_{k}$ with more than a specific threshold of certainty $\tau_{cer}$ denoted as
@@
When a single 2D mask spans multiple distinct objects that are spatially close in the 2D image but separated in 3D space (e.g., a vase in front of a couch), we refine it by applying DBSCAN clustering (Ester et al., 1996) to each $\hat{\mathcal{M}}_{t}^{(k)}$ to get the set of $\tilde{\mathcal{M}}_{t}^{(k)}$. This step separates the projected 3D points of $m$ into distinct clusters, each corresponding to a potential object. Each cluster is treated as a new candidate instance: we project it back to the image plane, generate the corresponding 2D mask. The final mask set for frame $t$ is then
$\mathcal{M}_{t}=\bigcup_{k=1}^{K}\tilde{\mathcal{M}}_{t}^{(k)}.$
@@ 3.2 Context-Aware CLIP Embedding
Given the refined 2D masks $\mathcal{M}_{t}$, we next compute semantic embeddings for each object candidate. A direct approach would be to crop the mask region and feed it into CLIP (Radford et al., 2021) or passing a raw 2D mask to OvSeg which is fine tuned on raw masks datasets. However, CLIP relies heavily on visual context, and isolated object crops often lead to ambiguous or incorrect embeddings. However, OvSeg struggles to overcome this issue, but in some scenarios, even an accurate mask does not contain meaningful semantics as shown in Fig. 2. To mitigate this, we construct a set of complementary visual crops for each mask that balance object detail with surrounding scene context.
@@
and re-apply our multi-crop CLIP embedding procedure. This ensures that each physically distinct object obtains its own semantic descriptor, even if they were originally fused into a single 2D segmentation.
After applying merging, we obtain the final set of refined 3D masks:
@@ 3.4 Object Retrieval
where $m$ denotes the name and attributes of the main object , $\mathcal{R}$ is a set of referenced object names and $\Omega$ encodes orientation constraints (e.g., “front of the cabinet”).A lightweight LLM extractor produces $\Pi(q)$ deterministically; for the full prompt, see C.1.
Candidate mining. For each object name $x\in\{m\}\cup\mathcal{R}$, we compute CLIP similarity using pre-computed object embeddings and retain the top-$K$ matches, where $K$ is either a fixed hyperparameter or an upper bound predicted by an LLM based on the object category and scene context. Each match corresponds to a 3D point cluster. We deliberately avoid thresholding CLIP scores and instead always select the top-$K$, since a single similarity threshold that works well for one query can be unreliable across scenes with substantially different similarity distributions. To remove duplicates, we voxelize all candidate clusters and discard any cluster whose voxelized support overlaps a larger neighbor beyond a fixed ratio.
View selection and VLM verification. Each candidate cluster is then projected into the set of RGB frames. We select the frame that maximizes 3D–2D overlap while penalizing occlusion by other candidate clusters. A VLM is prompted with the bounding box region and asked a binary question to determine whether the object is present. Only candidates passing this check are retained; for the prompt used in this verification step, see  C.2.
Orientation grounding. If $\Omega$ specifies an orientation, we collect canonical views of each candidate at discretized yaw bins. These views are tiled into a numbered grid, and a VLM is asked to select the index corresponding to the orientation token (e.g., front). The chosen index is mapped back to a yaw angle and stored with the candidate; for the prompt used in this step, see C.3.
Final reasoning. The remaining candidates for $m$, together with centroids of related objects and any orientation cues, are passed to an LLM. The LLM receives the original query and structured scene geometry and outputs the index of the final prediction; for the prompt used in this step, see C.4.
@@ 4 Experiments
We evaluate our framework on two tasks: (i) 3D open-vocabulary semantic segmentation, where the goal is to assign category labels to 3D object instances without task-specific training, and (ii) natural-language object retrieval, where the goal is to localize objects in a 3D scene given free-form text queries that may contain relational and orientation constraints. All experiments are conducted on a workstation equipped with a single NVIDIA RTX4090, with the exception of the VLM and LLM components, which are accessed through external APIs.
@@
Datasets. We conduct experiments on two standard benchmarks: Replica (Straub et al., 2019) and ScanNet (Dai et al., 2017). Replica provides high-quality synthetic RGB-D scans of indoor environments with accurate ground-truth meshes and semantic annotations, while ScanNet consists of large-scale real-world RGB-D sequences with manually annotated 3D semantic and instance labels. Following prior work, we use eight Replica scenes: room0, room1, room2, office0, office1, office2, office3, and office4, and eight ScanNet scenes: 0011_00, 0030_00, 0046_00, 0086_00, 0222_00, 0378_00, 0389_00, and 0435_00. This subset selection ensures comparability with previous zero-shot methods. For text–image alignment, we use the Eva02-L CLIP (Fang et al., 2023; Yang et al., 2023) vision-language model.
Baselines. We compare against both training-based and zero-shot 3D open-vocabulary segmentation methods. Training-based baselines include LERF (Kerr et al., 2023), OpenNeRF (Engelmann et al., 2024), LangSplat (Qin et al., 2024), and OpenGaussian (Wu et al., 2024). Zero-shot baselines include ConceptFusion (Jatavallabhula et al., 2023), ConceptGraphs (Gu et al., 2024), BBQ-CLIP (Linok et al., 2025), OpenMask3D (Takmaz et al., 2023), and HOV-SG (Werby et al., 2024). Note that HOV-SG reports results on a different subset of ScanNet, so we omit its numbers for fair comparison.
Results. Table 1 reports quantitative comparisons on Replica and ScanNet. Our method achieves the best performance across all evaluated approaches—including both training-based and zero-shot methods. In particular, the improvements in mIoU and fwIoU highlight the benefit of our context-aware embeddings and multi-view 3D refinement, which produce more consistent object representations than 2D-based pipelines. We also provide qualitative examples in Fig. 4. Visual comparisons on Replica scenes illustrate that our method yields more accurate segmentation boundaries and finer recognition of challenging categories. Our method consistently detects objects across categories with higher fidelity than competing approaches.
@@
Training-based methods
LERF 0.26 0.11
OpenNeRF 0.32 0.20
LangSplat 0.09 0.04
OpenGaussian 0.42 0.25
Zero-shot methods
@@ 4.2 Natural-Language Object Retrieval
5 Ablation studies
In this section, we assess the effectiveness of each part of the 3D semantic segmentation pipeline and also of 3D object retrieval. First of all, we have evaluated our progressive multi granularity level mask generation strategy. We compare our approach with Vanilla SAM model and also with different single granularity levels of Semantic-SAM model.
Table 3: Granularity-based mask generation impact
Methods Replica ScanNet
[…]
SAM 0.32 0.22 0.37 0.43 0.27 0.42
Semantic-SAM(Granularity Level=2) 0.34 0.25 0.51 0.54 0.32 0.43
Semantic-SAM(Granularity Level=4) 0.38 0.24 0.43 0.43 0.26 0.46
Semantic-SAM(Granularity Level=6) 0.26 0.11 0.14 0.17 0.09 0.16
Ours 0.38 0.29 0.56 0.61 0.36 0.46
To evaluate the effectiveness of our context-aware CLIP embedding generation approach, compare it to passing the raw masks to the OvSeg (Liang et al., 2023) model which is the finetuned version of the CLIP model on the raw masks dataset.
Table 4: Context-aware CLIP embedding impact
Methods Replica ScanNet
[…]
OvSeg 0.21 0.11 0.42 0.27 0.16 0.21
Context-Aware CLIP 0.38 0.29 0.56 0.61 0.36 0.46
For our approach, we have used Large version of CLIP model which has the same number of parameters as the finetuned OvSeg model. Our approach strongly enhances the CLIP model’s embeddings quality without any finetuning.
We have also evaluated our extension mechanism. Instead of relying on target mask size for extending the view around the target mask, we can leverage neighbor masks sizes to determine extension ratio. We define the large extension ration as the ratio that covers all the neighboring masks and the Huge as the ratio that covers all the neighbors of neighboring masks.
Table 5: Extension mechanism impact.
Methods Replica ScanNet
[…]
Neighbor coverage 0.34 0.28 0.45 0.51 0.28 0.36
Mask size 0.38 0.29 0.56 0.61 0.36 0.46
@@ 6 Conclusion
Appendix A 3D semantic segmentation
A.1 Labeling protocol
For open-vocabulary labeling, we use the standard CLIP paradigm (Radford et al., 2021): each refined 3D object mask is encoded into an embedding and compared against a set of text embeddings corresponding to candidate categories. The category with the highest similarity is assigned as the predicted label. This procedure enables zero-shot semantic segmentation without task-specific training, while our multi-view refinement makes the assignments more robust to occlusions and clutter.
A.2 Evaluation protocol
We follow the evaluation protocol in prior work. For each predicted point cloud instance, we assign a semantic label by finding the nearest ground-truth instance (based on centroid distance) and transferring its label. For each scene, we restrict the text prompts to the classes present in the ground-truth annotations, formatted as “a photo of <class name>.” We report mean accuracy (mAcc), mean intersection-over-union (mIoU), and frequency-weighted mIoU (fmIoU).
Appendix B 3D Object Retrieval
B.1 Implementation details
We use the EVA02 CLIP backbone (Fang et al., 2023) for visual–text alignment, and qwen2.5-vl-32b-instruct as the vision–language model for multimodal encoding. Our first LLM, gpt-5-mini, is employed for object extraction, while the second LLM, openai-o4-mini, handles the final decision-making stage.
B.2 Evaluation protocol
Following prior work, we report grounding accuracy at two IoU thresholds: A@0.1 and A@0.25. A prediction is considered correct if the IoU between the predicted and ground-truth bounding box exceeds the threshold. Accuracy is reported overall as well as separately for the four difficulty subsets.
Appendix C LLM Prompts for Object Retrieval
In this work, we use both LLMs and VLMs to interpret natural-language queries and convert them into structured outputs for 3D object retrieval. In this section, we provide the prompts used in our system, covering query parsing, main-object extraction, and candidate selection.
C.1 Query Parsing
The retrieval pipeline begins by converting each natural-language instruction into a structured representation containing the main object, related objects, the spatial relation between them, and any orientation constraints. This is obtained using the following prompt:
Your task is to strictly extract structured information from the given text. Format rules (follow exactly): 1) Line 1 — MAIN OBJECT: write only one object (the target), optionally followed by comma-separated visual attributes (color, shape, size, texture, material). No non-visual or spatial attributes. 2) Line 2 — RELATED OBJECTS: list all related objects with their visual attributes using: object_name[, attr1, attr2…]; object_name2[, attr1…] Do not repeat the main object. No spatial attributes. Leave blank if none. 3) Line 3 — RELATION KEYWORD: write exactly one of: closest, farthest, left, right, front, back, below, supported-by, above, supporting, between 4) Line 4 — ORIENTATION IMPORTANCE: list objects with required orientation in the form: object_name: orientation Allowed: front, back, left, right. Leave blank if none. Do not add placeholders, extra text, labels, punctuation, or formatting. Text: <user_instruction> Return exactly four lines.
This prompt provides a deterministic mapping from free-form language to a structured form suitable for geometric reasoning.
Although the prompt requests a spatial-relation keyword on Line 3, this information is not used by our retrieval pipeline. It is included only for analysis, and can be removed without affecting the retrieval system itself.
C.2 Object Presence Determination
After selecting the appropriate frame for each candidate, a red bounding box is drawn around its projected 2D region, and the annotated frame is provided to the VLM together with the following prompt:
Focus only on the red bounding box in the image. Ignore everything outside the box. Determine if the object <object_name> is present inside the box. It is valid even if only part of the object appears, as long as the main subject is included. Answer strictly with ’Yes’ or ’No’.
This prompt ensures that the VLM evaluates the candidate purely based on the content of the region indicated by the bounding box, while visually ignoring the rest of the frame.
C.3 Viewpoint Identification
For instructions that include orientation-dependent relations (such as “front of,” “left of,” or “to the right when facing the cabinet”), the system must determine which viewpoint of the candidate object corresponds to the requested orientation. To do this, we render viewpoints at sufficiently separated yaw angles (e.g., 90° increments), determined by camera locations relative to the object. These views are arranged into a single grid image, with each sub-image labeled by an index and containing a red bounding box marking the target object.
This grid image is then given to the VLM, which selects the sub-image matching the orientation specified in the instruction. The prompt used is:
In the provided image, multiple sub-images are arranged in a grid, each bordered in blue and labeled with a number at the top. In every sub-image, the target object ’<object_name>’ is indicated with a red bounding box. Ignore all other content. Determine which sub-image shows the object from the ’<orientation>’ view. Output only the index number above that sub-image.
This allows the VLM to choose the correct viewpoint among all candidates presented within the same image grid.
C.4 Candidate Selection and Final Retrieval
Once presence verification and viewpoint identification (when required) are completed, the system must choose the correct 3D instance of the main object. To do so, an LLM is provided with the original instruction together with the 3D centroids of all candidate object clusters. Each centroid corresponds to the geometric center of the candidate’s projected 3D point cloud. The LLM receives these centroids in plain text form along with the positions of referenced objects and, when applicable, the estimated orientation angles.
Depending on whether the instruction requires orientation reasoning, one of the two prompts below is used.
Without orientation.
If the instruction does not involve orientation-dependent relations, the LLM receives the following prompt:
Text: <user_instruction> The main object is <main>. It has the following candidate positions with indices (x, y, z): Index 0: (x=<x0>, y=<y0>, z=<z0>) Index 1: (x=<x1>, y=<y1>, z=<z1>) Index 2: (x=<x2>, y=<y2>, z=<z2>) Other objects and their positions are: <obj>[<k>]: (x=<x>, y=<y>, z=<z>) Instruction: Based on the text above, decide which candidate index of <main> is the correct target. Return only the numeric index (0, 1, 2, …). Do not return anything else.
Here, all coordinate placeholders (e.g., <x0>, <y>) are filled with the centroids of the 3D clusters, and name placeholders (e.g., <main>, <obj>) are filled with the corresponding object identifiers.
With orientation.
If the instruction requires orientation-dependent reasoning, the LLM receives the same information plus explicit orientation angles and their derived directional axes. The exact prompt is:
Text: <user_instruction> The main object is <main>. It has the following candidate positions with indices (x, y, z where z is height): Index 0: (x=<x0>, y=<y0>, z=<z0>) Index 1: (x=<x1>, y=<y1>, z=<z1>) Other objects and their positions are: <obj>[<k>]: (x=<x>, y=<y>, z=<z>) Orientation information (object, label) and angles in degrees: <obj> (<lab>): [<ang>] When facing <lab> (<ang>°): Forward = <ang>° (<dirF>), Left = <left>° (<dirL>), Right = <right>° (<dirR>), Back = <back>° (<dirB>). Instruction: Based on the text above, decide which candidate index of <main> is the correct target. Return only the numeric index (0, 1, 2, …). Do not return anything else.
Here, coordinate placeholders (e.g., <x0>, <x>, <y>, <z>) are filled with 3D centroids, name placeholders (e.g., <main>, <obj>) are filled with object identifiers, and orientation placeholders (e.g., <lab>, <ang>, <dirF>, <left>, <right>, <back>) are filled with the yaw angle and its derived directional axes.
In both cases, the LLM selects the final target purely from this structured geometric information and outputs a single integer index.
Appendix D Mask Merging
We use weighted averaging for our mask merging part. The weight of each mask is determined by its 3D relative position from the camera, amount of semantic the mask contains and the number of pixels in the mask. The position weight is determined by this formula:
$w_{d}=\begin{cases}e^{-\alpha(1.5-d_{avg})}&\text{if }d_{avg}\leq 1.5,\\ 1.0&\text{if }2.5\geq d_{avg}\geq 1.5,\\ e^{-\alpha(d_{avg}-2.5)}&\text{if }d_{avg}<2.5\end{cases}$
The $d_{avg}$ is the average depth of pixels in the mask and $\alpha=0.5$.
To measure semantics of each mask, we measure the cosine similarity of each mask with a set of indoor classes texts. The set contains 200 indoor classes and is generated by an LLM. The semantic weight is defined as $w_{s}=\displaystyle\max_{s\in S}s$.
The $w_{p}$ is the number of pixels in the mask. The final weight is calculated by:
$w=w_{d}\times w_{s}\times w_{p}$
Appendix E Hyperparameter Sensitivity
Context-aware embedding weights.
For each 2D object mask we construct five CLIP image embeddings:
$E_{s},\;E_{l},\;E_{h},\;E_{\text{hide}},\;E_{\text{mask}}\in\mathbb{R}^{d},$
corresponding respectively to (i) a small crop tightly around the object, (ii) a larger crop including more local context, (iii) a “huge” crop covering the object and its wider surround, (iv) a crop where the object is removed and only its surroundings are visible, and (v) an object-only crop obtained by masking out the background. Given these base embeddings, our normalized context-aware representation for an object is
$E_{\text{ctx}}=E_{s}+\alpha_{h}\,E_{h}+\alpha_{l}\,E_{l}-\alpha_{o}\,E_{\text{hide}}+\alpha_{m}\,E_{\text{mask}},$
where the scalar coefficients $\alpha_{h},\alpha_{l},\alpha_{o},\alpha_{m}\geq 0$ control how much we emphasize wide-field context ($E_{h}$), medium-scale context ($E_{l}$), suppress misleading background-only evidence ($E_{\text{hide}}$), and sharpen the signal from the object-only view ($E_{\text{mask}}$).
Sensitivity setup.
In the following experiments we study the effect of these weights on 3D semantic segmentation performance. For each parameter $\alpha\in\{\alpha_{h},\alpha_{l},\alpha_{o},\alpha_{m}\}$, we scale it by a factor $\lambda\in\{0.5,0.75,1.0,1.25,1.5\}$ while keeping the remaining coefficients fixed at their default values, and measure mIoU, mAcc, and fmIoU.
[…]
Figure 5: Hyperparameter sensitivity on Replica (scene room0) for the four weighting parameters $\alpha_{h},\alpha_{l},\alpha_{o},\alpha_{m}$. Performance is reported in mIoU, mAcc, and fmIoU as the scaling factor is varied.
[…]
Figure 6: Hyperparameter sensitivity on ScanNet (scene scene0011_00) for the four weighting parameters $\alpha_{h},\alpha_{l},\alpha_{o},\alpha_{m}$. Performance is reported in mIoU, mAcc, and fmIoU as the scaling factor is varied.
Appendix F Visualization
We include the following visualizations to enable a more detailed inspection of the results across different Replica scenes, facilitate comparison across methods, and to illustrate typical failure cases observed in challenging scenarios.